OpenStudy (anonymous):

Algebra Help Plz :) 2. What is the solution to the system of equations y = 3x + 3 and y = 2x + 1? A) no solution B) (-1, 3) C) infinite number of solutions D) (-2, -3)

OpenStudy (anonymous):

@Daniel56k

OpenStudy (anonymous):

plug in y=3x+3 into y=2x+1

OpenStudy (mathmale):

Why not actually solve this system yourself and then check your answers against the given possible answers? Both your equations begin with 'y'. Set one equal to the other, which in effect eliminates y. Then, find x. Next, calculate y.

OpenStudy (anonymous):

then it will look like this 3x+3=2x+1

OpenStudy (anonymous):

after solving that plug in your answer into either y=3x+3 or y=2x+1

OpenStudy (anonymous):

@Daniel56k y = 3x + 3 y = 2x + 1 x = -2 and y = -3

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

follow the step i gave you

OpenStudy (anonymous):

finish it up by finding x and plug x into equation 1 or 2 to get y

OpenStudy (anonymous):

okay subtract both side by -2x-3

OpenStudy (anonymous):

and that is how you find x then plug x into y=2x+1

OpenStudy (anonymous):

okay 3x-2x=-3+1 x=-2 y=2(-2)+1 y=-3

OpenStudy (mathmale):

Is your final answer a point? If so, please write it in the form (x, y).

OpenStudy (anonymous):

yes